# OmniJet-: The first cross-task foundation model for particle physics

### v1

@@ Appendix C Generative model trained on single-jet data
The result of the jet training is shown in FigureĀ 12, of the jet training in FigureĀ 13. In the case, we see a good agreement between the reconstructed tokens and the generated events. However, it seems to be somewhat more difficult for the model to accurately model for jets, which is also mirrored for this quantity in the combined model (see FigureĀ 6).
It is interesting to compare the result of OmniJet- with that of a different generative model. EPiC-FMĀ Birk etĀ al. (2023) was the first generative model trained on the JetClass dataset, utilizing flow matching and operating without tokenization. The result of the comparison can be seen in FigureĀ 14. The plots show the JetClass data, the reconstructed JetClass token from this work, the EPiC-FM generated events, and the OmniJet- generated events. We use the more challenging class for comparison.
The ratio plots under the main plots show the generated events compared to their respective truths: direct JetClass for EPiC-FM and Reconstructed JetClass tokens for OmniJet-. Hence, the ratios show how well the respective generative models learn to replicate their training data. In general, we see that both models are doing well. OmniJet- has a somewhat higher discrepancy in the tails of all distributions except for constituent and the number of constituents. The most difficult distribution is the constituent , with bumps in the tail, which could also be seen in FigureĀ 13.

### v2

@@ Appendix C Generative model trained on single-jet data
The result of the jet training is shown in FigureĀ 16, of the jet training in FigureĀ 17. In the case, we see a good agreement between the reconstructed tokens and the generated events. However, it seems to be somewhat more difficult for the model to accurately model for jets, which is also mirrored for this quantity in the combined model (see FigureĀ 6).
It is interesting to compare the result of OmniJet- with that of a different generative model. EPiC-FMĀ Birk etĀ al. (2023) was the first generative model trained on the JetClass dataset, utilizing flow matching and operating without tokenization. The result of the comparison can be seen in FigureĀ 18. The plots show the JetClass data, the reconstructed JetClass token from this work, the EPiC-FM-kintop generated events, and the OmniJet- generated events. We use the more challenging class for comparison.
The ratio plots under the main plots show the generated events compared to their respective truths: direct JetClass for EPiC-FM-kintop and Reconstructed JetClass tokens for OmniJet-. Hence, the ratios show how well the respective generative models learn to replicate their training data. In general, we see that both models are doing well. OmniJet- has a somewhat higher discrepancy in the tails of all distributions except for constituent and the number of constituents. The most difficult distribution is the constituent , with bumps in the tail, which could also be seen in FigureĀ 17.